OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

y=-3x-7 I believe.

OpenStudy (anonymous):

y=3x-9

OpenStudy (anonymous):

Sorry. The 2nd one is right.

OpenStudy (anonymous):

Put it into point slope form. Y-y1=m(x-x1) y-6=-3(x+1) Solve it out. We know the slope is -3 because parallel lines have the same slope.

OpenStudy (anonymous):

wouldn't it be 1/3 ?

OpenStudy (anonymous):

Parallel lines have the same slope, perpendicular have opposite reciprocal slopes. You're thinking of perpendicular, but in this case the slope of a parallel line would be -3.

OpenStudy (anonymous):

ohhh okay. i got it.